OpenStudy (anonymous):

Two competing internet companies have recently restructured their payment plans. Quick Net requires a $10 activation fee and charges $55 per month for their basic plan. Speedy Net requires no activation fee and charges $57 per month for their basic plan. When will the two plans be equal in cost?

OpenStudy (anonymous):

and How can x + 8 = -x + 7 be set up as a system of equations? y = x + 8 y = -x + 7 y = x + 8 y = -3x + 7 y = x + 5 y = -2x + 11 y = 8 y = -3x + 7

OpenStudy (anonymous):

please please please helop me oout soooo tired

OpenStudy (anonymous):

Let x be the number of months required to get the two plans equal Cost of the first company after x months is 10 + 55 x Can you express the cost of the second company after x months in terms of x?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

What is it on term of x?

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

If it charges 57 a month, how much money would you pay this company after 2 months. The answer is 2(57)= 114 how much money would you pay this company after 3 months? Answer is 3(57)=171 how much money would you pay this company after x months?

OpenStudy (anonymous):

How much money would you pay this company after x months? Answer; x(57)=57 x

OpenStudy (anonymous):

Now, you equate the two costs and solve for x

OpenStudy (anonymous):

10 + 55 x = 57 x 10 = 57 x - 55 x 10 = 2 x x =2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw